Date: Thu, 17 Sep 1998 18:58:46 -0700 (PDT)
To: 2mass
Subject: IPAC 2MASS WG Meeting #165 Minutes
Cc: chas, stiening, bgreen
Mime-Version: 1.0
Content-Transfer-Encoding: 7bit
Content-MD5: n+J2zheeBSkD/bWtO6nVgA==

           IPAC 2MASS Working Group Meeting #165 Minutes

Attendees: G. Kopan, S. Van Dyk, R. Beck, T. Evans, R. Hurt, J. White,
           J. Fowler, H. McCallon, S. Wheelock, R. Tam, D. Kirkpatrick,
           D. Engler, W. Wheaton


1.) Project Update 
2.) Southern OPS Update
3.) New Action Items 

     Since R. Cutri was not present, G. Kopan ran the meeting.
1.) Project Update 

     D. Kirkpatrick provided an update on the northern facility.  The facility
is out of summer shutdown, as of Sunday.  R. Stiening has been changing the
collimation and running focus tests for three nights.  R. Beck and D. 
Kirkpatrick have been processing the images and analyzing the tests.  The system
is not yet optimized.  E. Howard (UMass) is flying to Mt. Hopkins to assist
R. Stiening, for several more nights of testing.

2.) Southern OPS Update

     Southern OPS have been run on 980319s.  Some minor problems have been
found; as G. Kopan reports, two bugs were found and fixed in cross-scan 
photometric correction code.  W. Wheaton reports that the PSFs appear to be OK;
he reports "nothing new and frightening."  H. McCallon reports that in a few
cases, the ACT residuals are wandering off, such that the position reconstruc-
tion is not tracking the ACT.  However, this is quite infrequent, and H. 
McCallon finds "no real show-stoppers."  T. Evans reports that MAPCOR is "making
sense."  MAPCOR seems to be using the right PSFs and lookup tables.  MAPCOR is 
issuing some warnings on aperture mag-profile mag dispersions, but T. Evans 
believes that this is data-driven by a galaxy cluster in the scans.  As she 
says, it is a "balancing act" for the rejection criteria, between throwing out 
stars and keeping galaxies.  This may be more an issue for v3.0, than v2.1.

     G. Kopan asked T. Evans about the status of the artifact tuning; T. Evans
reports that all tuning is done for the latest PSFs.  However, a new issue
(unrelated to southern OPS) is the by-hand removal of artifacts from the Sampler
catalog, especially for the Beta Peg-affected scans.  

     S. Wheelock reports that all is fine from the calibration end, except that
CALMON must be rerun for this night using a linear fit.

     H. McCallon discussed the loading of the new USNO-A catalog, which R. Beck
will do by stopping production briefly at a good breakpoint.  This will 
undoubtedly shift some star positions in cal scans, as a warning to R. Hurt and
D. Kirkpatrick, on the QA end.  Some positions may shift by as much as 1 arcsec.
For survey scans with a small number of ACT stars, reconstructing to the new
USNO-A catalog should be more successful and should give a better indication of
the "walk-off" effect for position reconstruction dispersions.

     On the QA end for 980319s, R. Hurt reported "nothing glaring."  T. Evans
reiterated that there is no way around a MAPCOR renormalization, but this will
be determined after more nights have been run.

     J. Fowler asked about how the data acquisition rate is running relative to
expectations (this question was directed relative to tape consumption).  G. 
Kopan indicated that the rate is pretty close to the predicted rate.  R. Beck
said that the original tape use prediction was 70 Gby per DLT, but is actually
more like 40 Gby.  He has resorted to gzipping large files, to buy tape space.

     Finally, W. Wheaton asked about the faint mystery source seen in a scan 
overlap region on 971027n, that had apparently moved between the scans by about
1 arcsec.  J. Fowler thought that MPCAT had not yet been run, to determine 
whether this might be a known asteroid.  

3.) New Action Items

%                                                      %
&                 ACTION ITEMS                         %
%                                                      %

     No new action items were issued by R. Cutri, in his absence, but analysis
of the second southern OPS night, 980321s, should continue.